An electric motor driven conveyor belt is required to take 2,200 lb of dirt, rocks, etc per minute from an excavation site and raise it 15 ft vertically. If all friction in the machinery is neglected, what HP of motor is required?


1
Expert's answer
2021-04-28T20:18:21-0400

The HP (or power) of the motor can be found from the energy required to raise this mass of load for 15 ft and the time it takes:


"P=\frac{E}{t}=\frac{mgh}{t}=17695.7\space\frac{\text{ft}·\text{lbf}}{\text s},"

or 33 HP.